A study claims tha.t on average, male high school students spend at least 3.39 hours a day playing video games with a standard deviation of 2.05 hours. A random sample of 30 teenagers were surveyed and the results showed that these teenagers spend 4.12 hours a day playing video games.



Solve for the z-score of the sample mean of 4.12 hours

1
Expert's answer
2022-05-17T14:45:26-0400

The z-score of the sample mean is


"z=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{4.12-3.39}{2.05/\sqrt{30}}\approx1.9504"